# Automated Knowledge Component Generation and Knowledge Tracing for Coding Problems

### v2

@@ Abstract
Knowledge components (KCs) mapped to problems help model student learning, tracking their mastery levels on fine-grained skills thereby facilitating personalized learning and feedback in online learning platforms. However, crafting and tagging KCs to problems, traditionally performed by human domain experts, is highly labor-intensive. We present a fully automated, LLM-based pipeline for KC generation and tagging for open-ended programming problems. We also develop an LLM-based knowledge tracing (KT) framework to leverage these LLM-generated KCs, which we refer to as KCGen-KT. We conduct extensive quantitative and qualitative evaluations on a real-world student code submission dataset. We find that KCGen-KT outperforms existing KT methods and human-written KCs on future student response prediction. We investigate the learning curves of generated KCs and show that LLM-generated KCs result in a better fit than human-written KCs under a cognitive model. We also conduct a human evaluation with course instructors to show that our pipeline generates reasonably accurate problem-KC mappings.
Automated Knowledge Component Generation and
Knowledge Tracing for Coding Problems
Zhangqi Duan1, Nigel Fernandez1, Arun Balajiee Lekshmi Narayanan2, Mohammad Hassany2, Rafaella Sampaio de Alencar2, Peter Brusilovsky2, Bita Akram3, Andrew Lan1, University of Massachusetts Amherst 1, University of Pittsburgh 2, North Carolina State University 3 zduan@cs.umass.edu
1 Introduction
@@
In student modeling, an important task is to map problems (or items or questions) to specific skills or concepts, referred to as knowledge components (KCs). KCs provide an invaluable resource to model student learning Bier et al. (2014), estimating their mastery levels Corbett and Anderson (1994) on fine-grained units of knowledge. Accurately estimating student mastery levels on KCs helps enable both 1) teacher feedback, by showing this information in teacher dashboards, and 2) adaptive and personalized learning in online learning platforms or intelligent tutoring systems Huang et al. (2020), by tailoring instructions and content sequencing according to student knowledge levels. Identifying fine-grained KCs students struggle Rivers et al. (2016) with also enables content designers to develop targeted instructional content and practice problems for students.
@@
Basic arithmetic operations Math (+/+-*/+ - ∗ /)
@@
Table 1: Example programming problem from the CodeWorkout dataset with a sample student solution code, comparing KCs generated by our KCGen-KT framework to human-written KCs.
KCs are typically crafted by human domain experts, who also tag problems with KCs that students need to master to solve the problem correctly. This process can be highly labor-intensive, prone to bias and errors, and may not be scalable. There exist solutions to automate parts of this process using Natural Language Processing (NLP) tools, usually employing classification algorithms Pardos and Dadu (2017), to tag KCs to problems, which relies on having a predefined set of KCs. Recent advances in Large Language Models (LLMs) have shown potential in developing automated approaches for KC identification in addition to tagging, in domains such as math Ozyurt et al. (2024) and science Moore et al. (2024). Automatically generating KCs is challenging since KCs need to satisfy various criteria including being relevant to problems, being specific enough to provide teacher and student support, and being generalizable across settings. Another important aspect is that they need to satisfy cognitive science principles, i.e., student error rates on a KC should decrease as they attempt it more times, according to the power law of practice Snoddy (1926).
@@
Unlike other domains, generating KCs for open-ended programming problems that are common in the domain of computer science education has unique challenges. Writing code is inherently non-linear, with complex interactions between programming concepts and skills, and requires students to construct functioning code from scratch. Moreover, a programming problem can often have multiple valid solutions using different strategies, which may cover different sets of KCs. Prior work Hosseini and Brusilovsky (2013) uses a Java parser to convert a solution program into an Abstract Syntax Tree (AST) and reports ontological concepts at the lowest level as KCs. LLMs, with their advanced programming and reasoning abilities, are yet to be tested for automated KC generation and tagging for programming problems. Other recent works de Alencar et al. (2025) use AST root nodes in student code submissions as KCs and show that these KCs lead to good learning curve fit under student models Pavlik et al. (2009). Due to spatial constraints, see Section A in the Appendix for a more detailed review of related work.
1.1 Contributions
@@
In this paper, we explore using LLMs to automatically generate KCs for open-ended programming problems. We also develop an LLM-based knowledge tracing (KT) framework to leverage these LLM-generated KCs, which we refer to as KCGen-KT [^1] https://github.com/umass-ml4ed/kcgen-kt. Our contributions are summarized as follows:
  1. 1.
@@
    We develop a fully automated, LLM-based pipeline for KC generation and tagging. We first select a diverse set of representative student code submissions to each problem and then prompt GPT-4o OpenAI (2024), an advanced, proprietary LLM, to identify KCs that are required to solve the problem. Then, to aggregate KCs across problems and de-duplicate similar ones, we cluster KCs on semantic similarity, followed by summarizing each cluster into a KC description. Finally, we automatically tag problems with KCs according to the clustering results. Table 1 shows an example problem with the set of LLM-generated KCs.
  2. 2.
@@
    We develop an LLM-based KT method to leverage the textual descriptions of the generated KCs for the KT task. Our method explicitly captures student mastery levels on each KC and is thus interpretable, while also predicting both the actual student code submission and its correctness.
  3. 3.
@@
    We conduct an extensive quantitative and qualitative evaluation on the CodeWorkout dataset (DataShop, 2021) that contains real-world student code submissions to open-ended programming problems. Results show that KCGen-KT outperforms existing KT methods and human-written KCs on predicting future student performance. We also investigate the learning curves for these KCs and show that LLM-generated KCs have a comparable level of fit to human-written KCs under the performance factor analysis model. We also conduct a human evaluation to show that the KC tagging accuracy of our pipeline is reasonably accurate to human instructors.
2 Methodology
@@
Figure 1: Illustration of our three-step automated KC generation and tagging pipeline.
2.1 Automated KC Generation
@@
For KC generation, we use GPT-4o (OpenAI, 2024), an advanced proprietary LLM with strong reasoning and programming abilities. Illustrated in Figure 1, we generate KCs for a programming problem following 3 key steps: 1) generating KCs associated with each problem and their descriptions separately through few-shot prompting, 2) cluster KCs across all problems, and 3) summarizing each cluster to obtain a finalized description of each KC. We detail these steps below.
@@
Initial KC Generation For each programming problem, we prompt GPT-4o in a chain-of-thought manner, to generate a list of KCs that capture the underlying skills or concepts necessary to solve the problem. We also include a few carefully constructed in-context examples in our prompt as few-shot demonstrations. We prompt GPT-4o to convert human-written topic tags, such as “If-else”, from the CodeWorkout dataset into more fine-granular natural language descriptions, and use them as the in-context examples. We instruct the model to reason step-by-step: to first identify why a particular KC is relevant to the problem, and then generate a clear textual description of the KC. To help the LLM better understand what is required to solve a problem, we include correct student submissions as examples in the prompt. Since programming problems can often be solved in multiple valid ways, we include diverse examples to ensure comprehensive coverage of relevant KCs. Therefore, we apply a clustering algorithm to the CodeBERT (Feng et al., 2020) embeddings of all correct student submissions and sample one per cluster, with the number of clusters controlling the diversity of examples. Empirically, we find that this approach yields fine-grained, function-level KC descriptions. See Appendix E for the exact prompt used for all steps in our KC generation pipeline.
@@
Clustering KCs and Controlling Abstraction Level The KCs generated for each problem are initially fine-grained, often describing specific function-level skills or concepts. To control the level of abstraction and obtain more generalizable KC descriptions, we first compute the Sentence-BERT (Reimers and Gurevych, 2019) embedding of the textual description of each KC, then apply Hierarchical Agglomerative Clustering (HAC) using cosine similarity as the distance function. By adjusting the number of clusters, we can flexibly merge semantically similar KCs into broader categories, effectively controlling the abstraction level of our KC descriptions. This clustering process enables us to move from detailed skill-level descriptions to higher-level conceptual groupings, aligning the KCs with different pedagogical or analytical goals depending on the downstream application.
@@
Labeling KC Clusters Finally, we label each KC cluster by prompting GPT-4o to generate a single, informative name that represents the cluster. We use a chain-of-thought prompt that guides the model to first reason whether any KC in the cluster can represent the entire group. If such a KC exists, it is selected as the cluster label; otherwise, the model is instructed to synthesize a concise description that captures the shared meaning of KCs in the cluster. This process yields a final set of generated KCs across problems at the desired level of abstraction. As a final step, we obtain problem-KC mappings, i.e., a Q-matrix Barnes (2005), by mapping each initially generated KC for each problem to its corresponding summarized cluster label.
2.2 Improving Knowledge Tracing via LLM-generated KCs
@@
We now detail KCGen-KT, a novel LLM-based KT method that exploits KC semantics and explicitly models student mastery levels on each KC.
@@
KT Problem Formulation For open-ended programming problems, we define each student response to a problem as $x_{t}\coloneq(p_{t},\{w^{i}_{t}\},c_{t},a_{t})$, where $p_{t}$ is the textual statement of the problem, $\{w^{i}_{t}\}$ are the KCs associated with the problem, $c_{t}$ is the student code submission, and $a_{t}$ is the correctness of the submission; in most existing KT methods, $a_{t}$ is treated as binary-valued (correct/incorrect). Therefore, our goal is to estimate a student’s mastery level of each KC given their past responses, $x_{0},\ldots,x_{t}$, and use this estimate to predict both 1) the overall binary-valued correctness $a_{t+1}\in\{0,1\}$ and 2) the open-ended code $c_{t+1}$ submitted by the student on their next attempted problem $p_{t+1}$. Following previous work Shi et al. (2022), $a_{t}=1$ if the student-submitted code passes all test cases associated with the problem, and $a_{t}=0$ otherwise.
@@
Figure 2: Overview of our KCGen-KT’s model with the Llama 3 LLM as the backbone. KCGen-KT leverages KC semantics, tracking student mastery levels on each KC, to predict both correctness and the student code submission.
@@
KCGen-KT KCGen-KT leverages the KCs associated with a problem in two ways: 1) by improving the problem representation using the semantic information of KCs, and 2) by improving the student representation by building an interpretable student profile modeling student mastery levels on KCs.
@@
Following TIKTOC Duan et al. (2025), we use an open-source LLM, Llama 3 Llama Team (2024), as the backbone to predict both the overall correctness and actual open-ended student code in a token-by-token manner, in a multi-task learning approach. KCGen-KT differs from OKT Liu et al. (2022) by leveraging the content of the KCs, and from Code-DKT Shi et al. (2022) by using text embedding methods to embed the textual problem statement.
@@
Student Knowledge on KCs For each student, at each timestep $t$, KCGen-KT updates the student’s $512$-dimensional knowledge state vector $h\in\mathcal{R}^{512}$, through a long short-term memory (LSTM) Hochreiter and Schmidhuber (1997) network as in DKT Piech et al. (2015), given by $h_{t}=\text{LSTM}(h_{t-1},p_{t},c_{t})$. This knowledge state $h_{t}$ is compressed into a $k$-dimensional mastery vector $m_{t}\in[0,1]^{k}$, where $k$ is the total number of KCs, through a linear layer with weights $W_{m}$ and bias $b_{m}$, followed by a sigmoid function to map the values of $m_{t}$ to be in the range of $[0,1]$, given by $m_{t}=\sigma(W_{m}h_{t}+b_{m})$. Each dimension $j$ of $m_{t}$ denotes a student’s mastery level on the $j$th unique KC, with larger values denoting higher mastery.

### v3

# Automated Knowledge Component Generation for Interpretable Knowledge Tracing in Coding Problems

@@ Abstract.
Knowledge components (KCs) mapped to problems help model student learning, tracking their mastery levels on fine-grained skills thereby facilitating personalized learning and feedback in online learning platforms. However, crafting and tagging KCs to problems, traditionally performed by human domain experts, is highly labor-intensive. We present an automated, LLM-based pipeline for KC generation and tagging for open-ended programming problems. We also develop an LLM-based knowledge tracing (KT) framework to leverage these LLM-generated KCs, which we refer to as KCGen-KT. We conduct extensive quantitative and qualitative evaluations on two real-world student code submission datasets in different programming languages. We find that KCGen-KT outperforms existing KT methods and human-written KCs on future student response prediction. We investigate the learning curves of generated KCs and show that LLM-generated KCs result in a better fit than human-written KCs under a cognitive model. We also conduct a human evaluation with course instructors to show that our pipeline generates reasonably accurate problem-KC mappings.
Computer Science Education, Large Language Models, Knowledge Components, Knowledge Tracing, Student Modeling
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXccs: Applied computing Educationccs: Computing methodologies Natural language processingccs: Social and professional topics Student assessment
1. Introduction
@@
In student modeling, an important task is to map problems (or items) to specific skills or concepts, referred to as knowledge components (KCs). KCs provide an invaluable resource to model student learning (Bier et al., 2014), estimating their mastery levels (Corbett and Anderson, 1994) on fine-grained units of knowledge. Accurately estimating student mastery levels on KCs helps enable both 1) teacher feedback, by showing this information in teacher dashboards, and 2) adaptive and personalized learning in online learning platforms or intelligent tutoring systems (Huang et al., 2020), by tailoring instructions and content sequencing according to student knowledge levels. Identifying fine-grained KCs students struggle (Rivers et al., 2016) with also enables content designers to develop targeted instructional content and practice problems for students. Therefore, KC mastery estimation is fundamental to student modeling for personalization and recommendation.
KCs are typically crafted by human domain experts, who also tag problems with KCs that students need to master to solve the problem correctly. This process can be highly labor-intensive, prone to bias and errors, and may not be scalable. There exist solutions to automate parts of this process using Natural Language Processing (NLP) tools, usually employing classification algorithms (Pardos and Dadu, 2017), to tag KCs to problems, which relies on having a predefined set of KCs. Recent advances in Large Language Models (LLMs) have shown potential in developing automated approaches for KC identification in addition to tagging, in domains such as math (Ozyurt et al., 2024) and science (Moore et al., 2024). Automatically generating KCs is challenging since KCs need to satisfy various criteria, including being relevant to problems, being specific enough to provide teacher and student support, and being generalizable across settings. Another important aspect is that they need to satisfy cognitive science principles, i.e., student error rates on a KC should decrease as they attempt it more times, according to the power law of practice (Snoddy, 1926).
Table 1. Example programming problem from the CodeWorkout dataset with a sample student solution code, comparing KCs generated by our KCGen-KT framework to human-written KCs.
@@
Basic arithmetic operations Math (+/+-*/)
@@
Unlike other domains, generating KCs for open-ended programming problems that are common in the domain of computer science education has unique challenges. Writing code is inherently non-linear, with complex interactions between programming concepts and skills, and requires students to construct functioning code from scratch. Moreover, a programming problem can have multiple valid solutions using different strategies, which may cover different sets of KCs. Prior work (Hosseini and Brusilovsky, 2013) uses a Java parser to convert a solution program into an Abstract Syntax Tree (AST) and reports ontological concepts at the lowest level as KCs. LLMs, with their advanced programming and reasoning abilities, are yet to be tested for automated KC generation and tagging for programming problems.
1.1. Contributions
@@
In this paper, we explore using LLMs to automatically generate KCs for open-ended programming problems. We also develop an LLM-based knowledge tracing (KT) framework to leverage these LLM-generated KCs, which we refer to as KCGen-KT [^1] We will make our code publicly available.. Our contributions are summarized as follows:
  1. (1)
@@
    We develop an automated, LLM-based pipeline for KC generation and tagging. We first select a diverse set of representative student code submissions to each problem and then prompt GPT-4o (OpenAI, 2024), an advanced LLM, to identify KCs that are required to solve the problem. Then, to aggregate KCs across problems and de-duplicate similar ones, we cluster KCs on semantic similarity, followed by summarizing each cluster into a KC description. Finally, we automatically tag problems with KCs according to the clustering results. Table 1 shows an example problem with the set of LLM-generated KCs.
  2. (2)
@@
    We develop an LLM-based KT method to leverage the textual descriptions of the KCs for both actual student code prediction and correctness prediction. Our method explicitly captures student mastery levels on each KC, enabling interpretability. We introduce a soft-token conversion mechanism to inject the mastery level to the LLM input space for gradient flow through models.
  3. (3)
@@
    We conduct extensive quantitative and qualitative evaluation on two datasets (DataShop, 2021; de Freitas et al., 2023) that contain real-world student code submissions to open-ended programming problems. Results show that KCGen-KT outperforms existing KT methods and human-written KCs on predicting future student performance. We also investigate the learning curves for these KCs and show that LLM-generated KCs have a comparable level of fit to human-written KCs under the performance factor analysis model. We also conduct a human evaluation to show that the KC tagging accuracy of our pipeline is reasonably accurate to human instructors.
2. Related Work
2.1. Knowledge Component Generation
Traditional methods for KC creation and tagging rely on human domain experts to identify the knowledge requirements for solving a problem (Bier et al., 2014), a highly time-consuming process. Recent work has proposed automated approaches for KC discovery and tagging, employing data-driven approaches including the Q-matrix method (Barnes, 2005). In programming, (Hosseini and Brusilovsky, 2013) uses a rule-based parser to obtain ASTs with KCs identified at their lowest ontological level,  (Rivers et al., 2016) defines KCs as nodes in an AST followed by a learning curve analysis to identify KCs that students struggle with in Python programming, (Hoq et al., 2024) uses an AST-based neural network to identify student misconceptions, (Shi et al., 2024) presents a deep learning approach for KC attribution, and (Shi et al., 2023, 2024) learn latent KCs, lacking textual descriptions, by training deep learning models on KT data enforced with priors from pedagogical theory. Recent advances in LLMs have inspired automated approaches for descriptive KC generation in dialogues (Scarlatos et al., 2025), math (Ozyurt et al., 2024), science (Moore et al., 2024), and learning curve studies (Fan et al., 2025). However, we’re among the first approaches to present an automated LLM-based pipeline for KC generation and tagging for open-ended programming problems.
Figure 1. Illustration of our three-step automated KC generation and tagging pipeline.:
2.2. Knowledge Tracing
KT has been extensively studied within the student modeling literature (Corbett and Anderson, 1994). The classic KT task aims to estimate a student’s mastery of KCs from their responses to past problems and use these estimates to predict their future performance. Classic Bayesian knowledge tracing methods (Pardos and Heffernan, 2010; Yudelson et al., 2013) use latent binary-valued variables to represent student KC mastery. With the widespread adoption of neural networks, multiple deep learning-based KT methods were proposed with limited interpretability since student knowledge is modeled as hidden states in these networks. Most of these methods use long short-term memory networks (Hochreiter and Schmidhuber, 1997) or variants (Piech et al., 2015; Shin et al., 2021; Liu et al., 2024), with other variants coupling them with memory augmentation (Zhang et al., 2017), graph neural networks (Yang et al., 2020), or attention networks (Ghosh et al., 2020; Pandey and Karypis, 2019; Sun et al., 2024). KT methods have been applied to many different educational domains, including programming (Hoq et al., 2023; Shi et al., 2022; Zhu et al., 2022). Recent work has attempted to leverage LLMs to develop generative KT methods predicting exact student responses to programming problems (Duan et al., 2025; Fernandez and Lan, 2024; Liu et al., 2022).  (Sun et al., 2025) uses LLM to automatically construct Q-matrices capturing fine-grained KC relationship in KT, while their work lacks comparison with code-specific KT baselines, and does not explore LLM-based KC generation. To the best of our knowledge, we are the first to present an LLM-based KT method for programming problems that leverages the textual content of KC descriptions, modeling interpretable student mastery levels on each KC, for improved KT performance.
3. Methodology
@@
3.1. Automated KC Generation
@@
For KC generation, we use GPT-4o (OpenAI, 2024), an advanced proprietary LLM with strong reasoning and programming abilities. Illustrated in Figure 1, we generate KCs for a programming problem following three key steps: 1) generating KCs associated with each problem and their descriptions separately through few-shot prompting, 2) clustering KCs across all problems, and 3) summarizing each cluster to obtain a finalized description of each KC. All steps are performed independently for each dataset to maintain domain consistency. We detail these steps below.
3.1.1. Initial KC Generation
@@
For each programming problem, we prompt GPT-4o in a chain-of-thought manner to generate a list of KCs that capture the underlying skills or concepts necessary to solve the problem. We also include a few carefully constructed in-context examples in our prompt as few-shot demonstrations. We prompt GPT-4o to convert human-written topic tags, such as “If-else”, from each dataset into more fine-granular natural language descriptions, and use them as the in-context examples. We instruct the model to reason step-by-step: to first identify why a particular KC is relevant to the problem, and then generate a clear textual description of the KC. To help the LLM better understand what is required to solve a problem, we include correct student submissions as examples in the prompt. Since programming problems can often be solved in multiple valid ways, we include diverse examples to ensure comprehensive coverage of relevant KCs. Therefore, we apply a clustering algorithm to the CodeBERT (Feng et al., 2020) embeddings of all correct student submissions and sample one per cluster, with the number of clusters controlling the diversity of examples. Empirically, we find that this approach yields fine-grained, function-level KC descriptions. See Appendix B for the exact prompt used for all steps in our KC generation pipeline.
3.1.2. Clustering KCs and Controlling Abstraction Level
@@
The KCs generated for each problem are initially fine-grained, often describing specific function-level skills or concepts. To control the level of abstraction and obtain more generalizable KC descriptions, we first compute the Sentence-BERT (Reimers and Gurevych, 2019) embedding of the textual description of each KC, then apply Hierarchical Agglomerative Clustering (HAC) using cosine similarity as the distance function. By adjusting the number of clusters, we can flexibly merge semantically similar KCs into broader categories, effectively controlling the abstraction level of our KC descriptions. This clustering process enables us to move from detailed skill-level descriptions to higher-level conceptual groupings, aligning the KCs with different pedagogical or analytical goals depending on the downstream application.
3.1.3. Labeling KC Clusters
@@
Finally, we label each KC cluster by prompting GPT-4o to generate a single, informative name that represents the cluster. We use a chain-of-thought prompt that guides the model to first reason whether any KC in the cluster can represent the entire group. If such a KC exists, it is selected as the cluster label; otherwise, the model is instructed to synthesize a concise description that captures the shared meaning of KCs in the cluster. This process yields the final set of generated KCs across problems at the desired level of abstraction. As a final step, we obtain problem-KC mappings, i.e., a Q-matrix (Barnes, 2005), by mapping each initially generated KC for each problem to its corresponding summarized cluster label.
3.2. Improving Knowledge Tracing via LLM-Generated KCs
@@
We now detail KCGen-KT, a novel LLM-based KT method that leverages KC semantics and explicitly models student mastery levels on each KC, for improved KT performance.
3.2.1. KT Problem Formulation
@@
For open-ended programming problems, we define each student response to a problem as $x_{t}\coloneq(p_{t},\{w^{i}_{t}\},c_{t},a_{t})$, where $p_{t}$ is the textual statement of the problem, $\{w^{i}_{t}\}$ are the KCs associated with the problem, $c_{t}$ is the student code submission, and $a_{t}$ is the correctness of the submission; in most existing KT methods, $a_{t}$ is treated as binary-valued (correct/incorrect). Therefore, our goal is to estimate a student’s mastery level of each KC given their past responses, $x_{0},\ldots,x_{t}$, and use this estimate to predict both 1) the overall binary-valued correctness $a_{t+1}\in\{0,1\}$ and 2) the open-ended code $c_{t+1}$ submitted by the student on their next attempted problem $p_{t+1}$. Following previous work (Shi et al., 2022), $a_{t}=1$ if the student-submitted code passes all test cases associated with the problem, and $a_{t}=0$ otherwise.
@@
Figure 2. Overview of our KCGen-KT’s model with the Llama 3 LLM as the backbone. KCGen-KT leverages KC semantics, tracking student mastery levels on each KC, to predict both correctness and the student code submission.:
3.2.2. KCGen-KT
@@
KCGen-KT leverages the KCs associated with a problem in two ways: 1) by improving the problem representation using the semantic information of KCs, and 2) by improving the student representation by building an interpretable student profile that models student mastery levels on KCs.
@@
Following TIKTOC (Duan et al., 2025), we use an open-source LLM, Llama 3 (Llama Team, 2024), as the backbone to predict both the overall correctness and actual open-ended student code in a token-by-token manner, in a multi-task learning approach. KCGen-KT differs from OKT (Liu et al., 2022) by leveraging the content of the KCs, and from Code-DKT (Shi et al., 2022) by using text embedding methods to embed the textual problem statement.
3.2.3. Student Knowledge on KCs
@@
For each student, at each timestep $t$, KCGen-KT updates the student’s $512$-dimensional knowledge state vector $h\in\mathcal{R}^{512}$, through a long short-term memory (LSTM) (Hochreiter and Schmidhuber, 1997) network as in DKT (Piech et al., 2015), given by $h_{t}=\text{LSTM}(h_{t-1},p_{t},c_{t})$. This knowledge state $h_{t}$ is compressed into a $k$-dimensional mastery vector $m_{t}\in[0,1]^{k}$, where $k$ is the total number of KCs, through a linear layer with weights $W_{m}$ and bias $b_{m}$, followed by a sigmoid function to map the values of $m_{t}$ to be in the range of $[0,1]$, given by $m_{t}=\sigma(W_{m}h_{t}+b_{m})$. Each dimension $j$ of $m_{t}$ denotes a student’s mastery level on the $j$th unique KC, with larger values denoting higher mastery.
3.2.4. Soft KC Mastery Token Conversion